# Deblurring Neural Radiance Fields with Event-driven Bundle Adjustment

### v1

@@ Abstract.
Neural Radiance Fields (NeRF) achieve impressive 3D representation learning and novel view synthesis results with high-quality multi-view images as input. However, motion blur in images often occurs in low-light and high-speed motion scenes, which significantly degrade the reconstruction quality of NeRF. Previous deblurring NeRF methods are struggling to estimate information during the exposure time, unable to accurately model the motion blur. In contrast, the bio-inspired event camera measuring intensity changes with high temporal resolution makes up this information deficiency. In this paper, we propose Event-driven Bundle Adjustment for Deblurring Neural Radiance Fields (EBAD-NeRF) to jointly optimize the learnable poses and NeRF parameters by leveraging the hybrid event-RGB data. An intensity-change-metric event loss and a photo-metric blur loss are introduced to strengthen the explicit modeling of camera motion blur. Experiment results on both synthetic data and real captured data demonstrate that EBAD-NeRF can obtain accurate camera poses during the exposure time and learn sharper 3D representations compared to prior works.
@@
ccs: Computing methodologies Shape representations
@@ 1. INTRODUCTION
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) achieve 3D implicit representation learning and photo-realistic novel view synthesis results with high-quality 2D images and precise camera poses as input. In low-light scenes, a camera often requires a longer exposure time to obtain an image with sufficient brightness (Zamir et al., 2021b). A hand-held camera or a moving camera may cause motion blur in the captured image. Moreover, a high-speed moving camera can also cause motion blur, even in bright scenes with short exposure time. The blurred images will cause NeRF to learn a blurry 3D implicit representation, resulting in degraded quality of the synthesized novel view image. Thus, reconstructing sharp NeRF from blurry images is a practical problem for 3D reconstruction in low-light scenes (Yao et al., 2024).
@@
Figure 1. Motivation of our proposed method. In a low-light scene or with a high-speed moving camera, motion blur usually occurs in the captured images. To reconstruct sharp NeRF from blurry images, Deblur-NeRF (Ma et al., 2022) uses a deformable sparse kernel to model the blur process. BAD-NeRF (Wang et al., 2023) linearly interpolates the camera poses and jointly learns the start and end poses of camera motion. However, these methods are unable to model complex motion blur and lack supervision information during exposure time with only a photo-metric loss. With the proposed event-driven bundle adjustment, our EBAD-NeRF can recover the accurate camera motion trajectory and consequently reconstruct sharper novel view images with sharper learned NeRF.
Motion blur is essentially caused by the camera pose change during exposure time. Recent deblurring NeRF works try to model the camera motion and optimize the estimated camera pose change with photo-metric loss (Ma et al., 2022; Wang et al., 2023; Qi et al., 2023). As shown in Figure 1, Deblur-NeRF (Ma et al., 2022) propose the deformable sparse kernel (DSK) module to simulate the change of rays from the camera. The camera motion trajectory of blur is implicitly learned in an additional multilayer perceptron (MLP), which is unstable when facing complex and severe motion blur. BAD-NeRF (Wang et al., 2023) uses linear interpolation to model the motion trajectory. The starting and ending poses of the trajectory are jointly optimized with the parameters of NeRF. This strategy works well for most motion blur, but a handheld camera is not necessarily shaking linearly under long exposures, and this could lead to a failure of reconstruction under the linear interpolation modeling. Although the previous works have tried different motion blur modeling methods for deblurring NeRF, relying solely on photo-metric loss cannot form effective constraints on the camera trajectory estimate. Because there is inherent information loss within the exposure time in the input blurred image, and the modeling is limited by this. The bio-inspired event camera can measure brightness changing asynchronously with high temporal resolution (Gallego et al., 2020), which can make up for the information loss of blurry images. Recently, E2NeRF (Qi et al., 2023) uses both event data and blurry images to model the blur in terms of image formation and achieved better results than the image-based deblurring works. However, camera poses of this work are obtained by pre-deblurred images and are not learnable, which limits its performance.
@@
In this paper, we use learnable poses to model the camera motion trajectory. Event data is introduced into the camera pose learning by conducting an additional intensity-change-metric loss to effectively optimize the changes in adjacent poses, Unlike previous image-based works with only blurry images as supervision, which lacks constraints within the camera pose changes, we are able to fit the camera poses to the real motion trajectory during the blurring process and achieve sharper NeRF reconstruction. Specifically, for each view, we first use multiple learnable camera poses to model the camera motion trajectory and transfer them into the SE(3) space (Belta and Kumar, 2002; Wang et al., 2023). During the training, we jointly optimize these poses and the parameters of the NeRF network. A photo-metric blur loss is also used to supervise the modeling of blurry image formation. We introduce an intensity-change-metric event loss to supervise pose changes during the exposure time. The image formation is additionally strengthened by the event loss. The framework can estimate an accurate camera trajectory corresponding to the input blurry image, and a sharper NeRF is learned. We extend 5 blender synthetic scenes in Deblur-NeRF with event data and use a DAVIS-346 event camera (Taverni et al., 2018) to capture spatial-temporal aligned Event-RGB (ERGB) data with ground truth. Experiments on both synthetic data and real data demonstrate that our method can learn sharper neural radiation fields and more accurate camera motion trajectories compared with image deblurring methods and deblurring NeRF methods. To summarize, we present the following contributions:
@@
2) An intensity-change-metric event loss and a photo-metric loss are presented to supervise the modeled motion blur in terms of both camera pose change and image formation conjugatively.
3) Experiments on both extended synthetic data and real captured data validate that our method achieves accurate motion trajectory estimation and high-quality 3D implicit reconstruction of the scene with severe blurry images and corresponding event data.
@@ 2. RELATED WORK
Method RGB Event Motion Blur Model
Deblur-NeRF - 2D Blur Kernal
DP-NeRF - 3D Blur Kernal
Sharp-NeRF - Grid-Based Blur Kernal
@@
E2NeRF Poses from Pre-deblurred Images
@@
2.1. Neural Radiance Fields
The neural radiation fields model achieves impressive novel view synthesis results and inspires a lot of subsequent research (Gao et al., 2022). For example, (Barron et al., 2021, 2022; Verbin et al., 2022; Xu et al., 2022) improve the quality of learned geometry and synthesized novel views. Other works try to improve the training and rendering speed such as FastNeRF (Garbin et al., 2021), MobileNeRF (Chen et al., 2023) and EfficientNeRF (Hu et al., 2022). PixelNeRF (Yu et al., 2021) and RegNeRF (Niemeyer et al., 2022) realize reconstructing high-quality NeRF with few input views. RawNeRF (Mildenhall et al., 2022) can recover a clean NeRF from noisy raw images. HDR-NeRF generates high dynamic novel view images with low dynamic images as input. Besides, nerf also requires accurate camera poses from COLMAP as input. BARF (Lin et al., 2021), L2G-NeRF, and Nope-NeRF (Bian et al., 2023) explore reconstructing NeRF when camera poses are not provided.
Blurry Images are often captured in low-light scenes or with a high-speed moving camera on drones or robots. Reconstructing sharp NeRF from blurry images also becomes a challenging problem. As shown in Table 1, Deblur-NeRF (Ma et al., 2022) proposes a deformable sparse kernel module to effectively model the blurring process. However, the blurring kernel is optimized based on 2D-pixel location independently. DP-NeRF (Lee et al., 2023) proposes a rigid blurring kernel and generates a 3D deformation field, which is constructed as the 3D rigid motion of the camera for each view. Sharp NeRF (Lee et al., 2024) proposed a learnable grid-based kernel to obtain sharp output from neural radiance fields. Inspired by BARF, BAD-NeRF (Wang et al., 2023) conducts linear interpolation for the camera poses during the exposure time and jointly optimizes the poses with bundle adjustment.
2.2. Image Deblurring
Traditional deblurring algorithms focus on finding the suitable blur kernel of a blurry image to recover a sharp image. Hand-crafted features or sparse priors are used to tackle this problem blur (Chan and Wong, 1998; Krishnan et al., 2011; Xu et al., 2013). Deep learning deblurring works directly learn an end-to-end mapping from blurry to sharp images (Tao et al., 2018a; Wieschollek et al., 2017; Zhang et al., 2020). MPRNet (Zamir et al., 2021a) and SRN (Tao et al., 2018b) achieve impressive single-image deblurring results. In addition to single-image deblurring, (Kim et al., 2016; Zhang and Carin, 2014; Zhang and Yang, 2015) proposes deblurring schemes for continuous video images, which used the characteristics of video images(such as temporal consistency) to better predict sharp images, but this also brought greater calculation amount. He et al. (He et al., 2016) weights a combination of input images to avoid explicit calculation of blur kernels and effectively improves processing speed. Other works have also tried implementing multiple image deblurring using convolutional neural network and Unet framework (Wieschollek et al., 2016; Su et al., 2017). Son et al.(Son et al., 2021) proposes an effective recurrent video deblurring network and achieves state-of-the-art performance. However, the information missing from blurry images during exposure time makes the robustness of the image-based method decreases in the face of various motion blur.
@@ 2.3. Event Camera
Event camera is based on a bio=inspired vision senor that measures brightness changing asynchronously (Gallego et al., 2020). The high temporal resolution and high dynamic visual data capturing paradigm makes event cameras ideal for flow estimation (Shiba et al., 2022; Hagenaars et al., 2021; Gehrig et al., 2021; Akolkar et al., 2020; Pan et al., 2020), feature detection and tracking(Xu et al., 2020; Zhang et al., 2021, 2022) and images deblurring (Pan et al., 2019; Lin et al., 2020; Shang et al., 2021; Xu et al., 2021). Pan et al. (Pan et al., 2019) propose a simple event-based double integral model (EDI) based on event generation to establish the connection between blurry images and events. Shang et al. (Shang et al., 2021), Jiang et al.(Jiang et al., 2020) and Sun et al. (Sun et al., 2022) using learning-based network to recover sharp images with events.
Recently, event-based NeRF works have emerged. Ev-NeRF (Hwang et al., 2023) models the measurement of the event sensor to learn grayscale neural radiance fields derived from the event stream. EventNeRF (Rudnev et al., 2023) reconstructs color NeRF with a color event camera. e-NeRF (Low and Lee, 2023) improves Ev-NeRF and EventNeRF for non-uniform camera motion. DE-NeRF (Ma et al., 2023) uses both RGB images and events to learn a deformable NeRF. E2NeRF first reconstructs a sharp NeRF with blurry images and corresponding event data and achieves state-of-the-art performance. However, it only explicitly models the blurring processing at the imaging aspect. For camera motion during the exposure time, E2NeRF relies on pre-deblurred images with EDI (Pan et al., 2019) algorithm and structure-from-motion method COLMAP for pose estimation. Notably, our proposed EBAD-NeRF transforms camera motion trajectories into a learnable model with event-driven bundle adjustment. The framework unlocks the potential of event information in explicit modeling of blur processing at the level of both image changes and camera motion.
Figure 2. Overview of our method. An image with motion blur is caused by the camera moving during the exposure time. We temporal evenly sample $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ and transfer them into SE(3) as learnable parameters. Then, the poses are input into the NeRF network to model the blurry image and event stream generation. The intensity-change-metric loss can optimize the modeling in terms of camera motion. The photo-metric loss can optimize and optimize the modeling in terms of image formation. With the joint learning of NeRF parameters $\theta$ and camera poses, a sharp NeRF is learned eventually.
@@ 3. METHOD
We introduce EBAD-NeRF to simultaneously learn the poses of the camera’s motion process and generate sharp neural radiance fields using both blurred images and the event stream within the corresponding exposure time. Events serve two key purposes within the framework: (1) optimizing camera trajectories within the exposure time, and (2) contributing to the physical formation of motion blur in the image by providing information on light intensity changes. Moreover, we employ a photo-metric blur loss to supervise blur formation in the RGB domain. Figure 2 shows the overview of our approach. Sec. 3.1 introduces the motion blur formation in the neural radiance fields. Sec. 3.2 illustrate the proposed event-driven bundle adjustment method. Sec. 3.3 shows the loss function and the derivation, and Sec. 3.4 describes the implementation details of our method.
@@ 3.1. Motion Blur Formation in Static 3D Scene
In a static 3D scene, camera motion blur is caused by the changes of camera pose $\mathbf{P}$ during the exposure time. A digital image $\mathcal{I}$ is obtained by measuring the integration of light intensity $I$ on the image sensor.
@@
where ${t_{start}}$ and ${t_{end}}$ is the start and end time of exposure. The light intensity $I$ is only corresponding to the camera pose, and we can express $\mathbf{P}$ as a function of time $\mathbf{P}=p(t)$. Then, a blurry image $\mathcal{B}$ can be expressed as:
@@
In NeRF (Mildenhall et al., 2021), the rendering results are closely related to camera poses. Given a camera pose $\mathbf{P}=P(t)$, for each pixel of the imaging plane, we can obtain a ray $\mathbf{r}$ that emits from the optical center of the camera and passes through the pixel $\mathbf{x}$. With stratified sampling, we can divide the part of this ray starting with $l_{near}$ and ending with $l_{far}$ into N equal parts and randomly sample points in each part. For each of the $N$ sampled points, we input the 3D coordination $\mathbf{o}$ in the world coordinate system and 2D view the direction $\mathbf{d}$ represented by the ray $\mathbf{r}$ into NeRF MLP $F_{\theta}$ with network parameters $\theta$:
(3) $(\mathbf{c},\sigma)=F_{\mathbf{\theta}}(\gamma_{o}(\mathbf{o}),\gamma_{d}(\mathbf{d})),$
the output of which is color $\mathbf{c}$ and density $\sigma$ and $\gamma(\cdot)$ is an encoder for a higher $K+1$ dimension of input of the network:
@@
we can get the final color value of the pixel $\mathbf{x}$ passed by the ray $\mathbf{r}$ corresponding to the pose $\mathbf{P}$. $\delta_{i}=l_{i+1}-l_{i}$ is the distance between adjacent sampled points, and $T_{i}$ is the transparency between $l_{n}$ and sampled point.
If we assume that the learned NeRF MLP is sharp, we can model the image motion blur by discretizing Eq. (2) with $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ sampled evenly over the exposure time:
(6) ^$\hat{\mathcal{B}}(\mathbf{x})=\frac{1}{p}\sum_{i=1}^{p}C(P(t),\mathbf{x}),\mathbf{x}\in\mathcal{X},\$
where $\mathcal{X}$ represents the pixels of the image sensor. Since we use virtual sharp image color $C(P(t),\mathbf{x})$ to replace the light intensity $I(P(t))$, we need to multiply each item by a weight of time in Eq. (6). As the poses are temporally evenly sampled, we can directly use an average $\frac{1}{p}$ to represent the weight.
At this point, we have established the connection between image motion blur, camera pose, and neural radiation field.

### v2

@@ Abstract.
Neural Radiance Fields (NeRF) achieves impressive 3D representation learning and novel view synthesis results with high-quality multi-view images as input. However, motion blur in images often occurs in low-light and high-speed motion scenes, which significantly degrades the reconstruction quality of NeRF. Previous deblurring NeRF methods struggle to estimate pose and lighting changes during the exposure time, making them unable to accurately model the motion blur. The bio-inspired event camera measuring intensity changes with high temporal resolution makes up this information deficiency. In this paper, we propose Event-driven Bundle Adjustment for Deblurring Neural Radiance Fields (EBAD-NeRF) to jointly optimize the learnable poses and NeRF parameters by leveraging the hybrid event-RGB data. An intensity-change-metric event loss and a photo-metric blur loss are introduced to strengthen the explicit modeling of camera motion blur. Experiments on both synthetic and real-captured data demonstrate that EBAD-NeRF can obtain accurate camera trajectory during the exposure time and learn a sharper 3D representations compared to prior works.
@@
copyright: noneccs: Computing methodologies Computer vision
@@ 1. Introduction
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) achieves 3D implicit representation learning and photo-realistic novel view synthesis results with high-quality 2D images and precise camera poses as input. In low-light scenes, a camera often requires a longer exposure time to obtain an image with sufficient brightness (Zamir et al., 2021b), and a handheld camera may cause motion blur in the captured image. Moreover, a high-speed moving camera can also cause motion blur, even in bright scenes with short exposure time. The blurred images will cause NeRF to learn a blurry 3D implicit representation, resulting in degraded quality of the synthesized novel view images. Thus, it is a practical problem to reconstruct a sharp NeRF from blurry images when facing low-light and high-speed scenes (Yao et al., 2024).
@@
Figure 1. Motivation of our proposed method. In a low-light scene or with a high-speed moving camera, motion blur usually occurs in the captured images. To reconstruct sharp NeRF from blurry images, Deblur-NeRF (Ma et al., 2022) uses a deformable sparse kernel to model the blur process. BAD-NeRF (Wang et al., 2023) linearly interpolates the camera poses and jointly learns the start and end poses of camera motion. However, these methods are unable to model complex motion blur and lack supervision information during exposure time with only a photo-metric loss. With the proposed event-driven bundle adjustment, our EBAD-NeRF can leverage the event to recover the accurate camera trajectory and learn a sharper NeRF, resulting in sharper novel view image rendering than the previous methods.
For static scenes, motion blur is caused by the camera pose change during the exposure time. The core problem of reconstructing a sharp NeRF is accurately restoring the motion trajectory to better model the motion blur formation. Correspondingly, we need to establish a connection between scene radiance fields and camera pose change. Recent deblurring NeRF works try to resolve this by using a learnable blur kernel (Ma et al., 2022; Lee et al., 2023, 2024) or interpolating the camera poses with learnable poses at the start and end of the exposure (Wang et al., 2023). However, they solely implicitly model the motion and are supervised by RGB images, which is insufficient to establish a strong connection between camera pose change and NeRF parameters. This results in unstable network training and a decrease in performance for complex and severe motion blur. As in Figure 1, Deblur-NeRF (Ma et al., 2022) and Bad-NeRF (Wang et al., 2023) all learn an inaccurate motion trajectory and an inferior NeRF with blurry rendering results.
The bio-inspired event camera can measure brightness change asynchronously with high temporal resolution (Gallego et al., 2020), which makes up for the information loss in blurry images. Recently, E2NeRF (Qi et al., 2023) uses the event data and blurry images to enhance the model of motion blur in terms of image formation and achieves better results than the image-based deblurring works. However, it uses pre-deblurred images and COLMAP (Schonberger and Frahm, 2016) to estimate camera poses, which is not learnable during training, and the pose change information in the event data is not fully modeled and utilized.
@@
In this paper, we model the motion blur in terms of both motion trajectory and camera imaging. Event data is introduced to establish an explicit connection between the actual camera motion trajectory and the target sharp NeRF by conducting an intensity-change-metric event loss. Specifically, for each view, we first use multiple learnable camera poses to model the camera motion trajectory and transfer them into the SE(3) space (Belta and Kumar, 2002; Wang et al., 2023). During the training, we jointly optimize these poses and the parameters of the NeRF network. We introduce an intensity-change-metric event loss to supervise pose changes during the exposure time. A photo-metric blur loss is also used to supervise the modeling of blurry image formation. The image formation is also additionally strengthened by the event loss. We extend 5 blender synthetic scenes in Deblur-NeRF with event data and use a DAVIS-346 event camera (Taverni et al., 2018) to capture spatial-temporal aligned Event-RGB (ERGB) real data with ground truth. Experiments on both synthetic and real data demonstrate that our method can learn superior NeRF and more accurate camera motion trajectories than previous image-based or ERGB-based deblurring NeRF methods. The rendering results are also better than the state-of-the-art image-based or ERGB-based image deblurring methods. To summarize, we present the following contributions:
@@
2) An intensity-change-metric event loss and a photo-metric blur loss are presented to supervise the modeled motion blur in terms of both camera pose change and image formation conjugatively.
3) Experiments on extended synthetic data and real-captured data validate that our method achieves accurate motion trajectory estimation and high-quality 3D implicit reconstruction of the scene with severe blurry images and corresponding event data.
@@ 2. Related Work
2.1. Neural Radiance Fields
The neural radiation fields model achieves impressive novel view synthesis results and inspires a lot of subsequent research (Gao et al., 2022). Some works (Barron et al., 2021, 2022; Verbin et al., 2022; Xu et al., 2022) improve the quality of learned geometry and synthesized novel views. Other works try to improve the training and rendering speed (Garbin et al., 2021; Chen et al., 2023; Hu et al., 2022). Besides, BARF (Lin et al., 2021), L2G-NeRF, and Nope-NeRF (Bian et al., 2023) explore reconstructing NeRF without camera poses estimated by COLMAP (Schonberger and Frahm, 2016) and input images.
Blurry Images are often captured in low-light scenes or with a high-speed moving camera on drones or robots. Reconstructing sharp NeRF from blurry images also becomes a challenging problem. As shown in Table 1, Deblur-NeRF (Ma et al., 2022) proposes a deformable sparse kernel module to effectively model the blurring process. However, the blurring kernel is optimized based on 2D-pixel location independently. DP-NeRF (Lee et al., 2023) proposes a rigid blurring kernel and generates a 3D deformation fields, which is constructed as the 3D rigid motion of the camera for each view. Sharp NeRF (Lee et al., 2024) proposed a learnable grid-based kernel to obtain sharp output from neural radiance fields. Inspired by BARF, BAD-NeRF (Wang et al., 2023) conducts linear interpolation for the camera poses during the exposure time and jointly optimizes the poses with bundle adjustment.
2.2. Image Deblurring
Traditional deblurring algorithms focus on finding the suitable blur kernel of a blurry image to recover a sharp image. Hand-crafted features or sparse priors are used to tackle this problem (Chan and Wong, 1998; Krishnan et al., 2011; Xu et al., 2013). Deep learning deblurring works directly learn an end-to-end mapping from blurry images to sharp images (Tao et al., 2018a; Wieschollek et al., 2017; Zhang et al., 2020). MPRNet (Zamir et al., 2021a) and SRN (Tao et al., 2018b) achieve impressive single-image deblurring results. However, the information missing from blurry images during exposure time decreases the robustness of the image-based methods and limits their performance when facing various and severe motion blur in low-light or high-speed scenes.
@@ 2.2. Image Deblurring
Method RGB Event Motion Blur Trajectory Modeling
Deblur-NeRF - 2D Blur kernel
DP-NeRF - 3D Blur kernel
Sharp-NeRF - Grid-based Blur kernel
@@
E2NeRF Fixed Poses from Pre-deblurred Images
@@ 2.3. Event Camera
The event camera is based on a bio-inspired vision sensor that measures brightness changing asynchronously (Gallego et al., 2020). The high temporal resolution and high dynamic visual data capturing paradigm makes it ideal for flow estimation (Shiba et al., 2022; Hagenaars et al., 2021; Gehrig et al., 2021; Akolkar et al., 2020; Pan et al., 2020), feature detection and tracking (Xu et al., 2020; Zhang et al., 2021, 2022) and images deblurring (Pan et al., 2019; Lin et al., 2020; Shang et al., 2021; Xu et al., 2021). Pan et al. (Pan et al., 2019) propose a simple event-based double integral model (EDI) based on event generation to establish the connection between blurry images and events. Other works (Shang et al., 2021; Jiang et al., 2020; Sun et al., 2022) use learning-based networks to recover sharp images with events.
Recently, some event-based NeRF works have emerged. Ev-NeRF (Hwang et al., 2023) models the measurement of the event sensor to learn grayscale neural radiance fields derived from the event stream. EventNeRF (Rudnev et al., 2023) reconstructs color NeRF with a color event camera. e-NeRF (Low and Lee, 2023) improves Ev-NeRF and EventNeRF for non-uniform camera motion. SpikeNeRF (Zhu et al., 2024) derives a NeRF-based volumetric scene representation from spike camera data. DE-NeRF (Ma et al., 2023) uses RGB images and events to learn a deformable NeRF. E2NeRF is the first to reconstruct a sharp NeRF with blurry images and corresponding event data and achieves state-of-the-art performance. However, it only explicitly models the blurring processing at the imaging aspect. For camera motion trajectory during the exposure time, E2NeRF uses EDI (Pan et al., 2019) to pre-deblurred images and COLMAP to estimate the camera poses that can not be optimized during training. Ev-DeblurNeRF (Cannici and Scaramuzza, 2024) extends from DP-NeRF (Lee et al., 2023) with a novel EDI loss and an eCRF network to simulate the generation of events more realistically and achieves a remarkable deblurring effect. It also uses extra continuous events between spare image views to enhance the NeRF learning, which differs from the setting of E2NeRF and ours.
Figure 2. Overview of our method. For each view of the static scene, an image with motion blur is caused by the camera moving during the exposure time. We temporal evenly sample $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ on the motion trajectory and transfer them into SE(3) as learnable variables $\{\mathbf{T}_{i}\}_{i=1}^{p}$. With the poses, we model the blurry image and event generation in our network and jointly optimize the NeRF parameters $\theta$ and motion trajectory with the ERGB data to learn a sharp NeRF eventually.
@@ 3. Method
We introduce EBAD-NeRF to simultaneously learn the poses during the camera motion blur process and sharp neural radiance fields with blurred images and the event stream within the corresponding exposure time. Events serve two key purposes within the framework: (1) optimizing camera trajectories within the exposure time, and (2) contributing to the physical formation of motion blur in the image by providing information on light intensity changes. Moreover, we employ a photo-metric loss to supervise blur formation in the RGB domain. Figure 2 is the overview of our approach.
@@ 3.1. Motion Blur Formation in Static 3D Scene
In a static 3D scene, camera motion blur is caused by the changes of camera pose $\mathbf{P}$ during the exposure time. A digital image $\mathcal{I}$ is obtained by measuring the integration of light intensity $I$ with respect to time $t$ on the image sensor:
@@
where ${t_{start}}$ and ${t_{end}}$ are the start and end time of the exposure period. Since light intensity $I$ is only corresponding to the camera pose in the static scene, we can express $\mathbf{P}$ as a function of time $\mathbf{P}=p(t)$ and a blurry image $\mathcal{B}$ can be expressed as:
@@
In NeRF (Mildenhall et al., 2021), given a camera pose $\mathbf{P}=P(t)$, for each pixel of the imaging plane, we can obtain a ray $\mathbf{r}$ that emits from the optical center of the camera and passes through the pixel $\mathbf{x}=(x,y)$. With stratified sampling, we can divide the part of this ray starting with $l_{near}$ and ending with $l_{far}$ into N equal parts and randomly sample one point in each part. For each of the $N$ sampled points, we input its 3D coordination $\mathbf{o}$ in the world coordinate system and the 2D view direction $\mathbf{d}$ represented by the ray $\mathbf{r}$ into NeRF MLP $F_{\theta}$ with network parameters $\theta$:
(3) $(\mathbf{c},\sigma)=F_{\mathbf{\theta}}(\gamma_{o}(\mathbf{o}),\gamma_{d}(\mathbf{d})).$
The outputs are color $\mathbf{c}$ and density $\sigma$ of the point, and $\gamma(\cdot)$ encodes the input to a higher $K+1$ dimension:
@@
we can get the final color value $C(\mathbf{r},\mathbf{x})=C(\mathbf{P},\mathbf{x})$ of the pixel $\mathbf{x}$ passed by the ray $\mathbf{r}$ corresponding to the pose $\mathbf{P}$. $\delta_{i}=l_{i+1}-l_{i}$ is the distance between adjacent sampled points, and $T_{i}$ is the transparency between $l_{near}$ and the sampled point.
If we assume that the learned NeRF MLP is sharp, we can model the image motion blur by discretizing Eq. (2) with $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}=\{P(t_{i})\}_{i=1}^{p}$ sampled evenly over the exposure time $t_{start}$ to $t_{end}$:
(6) ^$\hat{\mathcal{B}}(\mathbf{x})=\frac{1}{p}\sum_{i=1}^{p}C(P(t_{i}),\mathbf{x}),\mathbf{x}\in\mathcal{X},\$
where $\mathcal{X}$ represents the pixels of the image sensor. Since we use virtual sharp image color $C(P(t_{i}),\mathbf{x})$ to replace the light intensity $I(P(t_{i}))$ of Eq. (2), we need to multiply each item by a weight of time. As the poses are temporally evenly sampled, we can directly use an average as $\frac{1}{p}$ to represent the weight as in Eq. (6).
At this point, we establish a connection between image motion blur, camera pose, and neural radiation fields.